# Morse-Bott inequalities, Topology Change and Cobordisms to Nothing

### v1

@@ 5 Stories with more than one ending
In the above sections, only a single BoN/EotW brane was considered for each case. However, instances of multi-centered BoN solutions eventually colliding or networks of EotW branes intersecting are the natural generalization. These cases have been previously studied in the literature (see Horowitz:2002cx ; Freivogel:2007fx ; Kleban:2011pg ; BoNColTBA and Angius:2023uqk ; Angius:2024zjv ; GarciaEtxebarria:2024jfv , respectively), and as we will see, we can learn important qualitative properties through our approach.
@@ 5.1 Seeing double: Collisions of BoN’s
As explained in section 2.1, bubbles of nothing have some non-zero probability of nucleating per unit volume. It is then expected, that, in a large macroscopic region and for a long enough period of time, several BoN nucleate and start growing, eventually approaching each other’s surfaces. Considering our lower-dimensional theory to result from compactification on a compact surfaces $\mathcal{C}_{n}$, a bubble solutions will be associated to a bordism and Morse function choice, $(\mathcal{B}_{n+1},\rho)$. Two different BoN can on principle be associated to $\mathcal{B}_{n+1}$ and $\mathcal{B}_{n+1}^{\prime}$, such that $\mathcal{C}_{n}=\partial\mathcal{B}_{n+1}=\partial\mathcal{B}_{n+1}^{\prime}$, without necessarily the two bordisms being homeomorphic. Moving then radially from the surface of one BoN to the other, this can be seen as traveling along some compact $(n+1)$-manifold $\mathcal{X}_{n+1}$ given by
@@
As the surfaces of the two bubbles approach, the volume of $\mathcal{X}_{n+1}$ shrinks. Whilst the eventual collision of the two tips can reach Planckian curvatures, as well as being affected by the backreaction of large matter densities being dragged by the bubble, one can also have these problems well before the BoN surface if $\mathcal{X}_{n}$ cannot be contracted to a point in a single topology change, with the curvatures of internal cycles of $\mathcal{X}_{n+1}$ blowing up before that of the whole manifold. Having a single topology change translates in the Morse function interpolating between the tips of the two bubbles having only critical two points, respectively with index 00 and $n+1$, associated with the tip of each BoN. From (3.14), this translates in
@@
which implies
@@
In the absence of torsion (i.e., all $t_{k}(\mathcal{M})\equiv 0$), a homology sphere solution $\vec{b}(\mathcal{X}_{n+1})=\vec{b}(\mathbb{S}^{n+1})$ is compatible [^22] Note that this does not imply such solution exists, but rather that there is no immediate obstruction to its existence. with the above (5.2) and (5.4) equations, by choosing [^23] For this choice, the resulting ${\mathrm{D}}\mathcal{B}_{n+1}\simeq\left(\mathcal{B}_{n+1}\cup\overline{\mathcal{B}_{n+1}}\right)/_{\sim_{\partial\mathcal{B}_{n+1}}}$ is called double of $\mathcal{B}_{n+1}$. $\mathcal{B}_{n+1}^{\prime}\simeq\mathcal{B}_{n+1}$ and saturating (3.8). This is illustrated for $\mathcal{C}_{2}\simeq\Sigma_{2}$ in Figure 10. This means that on principle there should not be a topological obstruction to the collision of two BoN with the $\mathcal{X}_{n+1}$ compact manifold between them shrinking to a point.
@@
As a final comment, we note that, unlike in the case of a single BoN, which required the presence of D$(8-p)$-branes as cobordisms defects when considering $p$-form fluxes, here their presence is much more mild than the intermediate topology changes commented above. D$(8-p)$-branes, located along along $\mathcal{B}_{n+1}$ and $\mathcal{B}^{\prime}_{n+1}$, needed to kill the $p$-form charges, have opposite orientations, and thus will annihilate each other when they collide as $\mathcal{X}_{n+1}$ closes. Another way to see this is by considering that the (co)homology groups of $\mathcal{X}_{n+1}$ and $\mathcal{C}_{n}$ are different, and as commented above, $H_{k}(\mathcal{X}_{n+1},\mathbb{Z})\simeq H^{k}(\mathcal{X}_{n+1},\mathbb{Z})$ can be trivial for $k=1,\,\dots,\,n$ in the absence of torsion, meaning that the number of $p$-form fields/charges is no longer conserved on the compact manifold connecting the two BoN.
[…]
@@ 5.2 Going a bordism further: Intersection of EotW branes
We finally consider the cases with more than one bordism to nothing, in the language of EotW branes. Intersection of these was studied in Angius:2023uqk ; Angius:2024zjv ; GarciaEtxebarria:2024jfv , and in this section we will study them from the point of view of topology changes between “cobordant” bordisms to nothing.
From the EFT point of view, the single EotW brane from section 2.2 was generalized to $k$ intersecting EotW branes in Angius:2023uqk . For this the following lower-$d$ EFT action is considered,
@@
where the moduli space metric has unit diagonal, $\mathsf{G}_{ii}=1$, as well as the following ansatz for a conformally flat metric (see Angius:2024zjv for EotW solutions with non-conformally flat backgrounds):
@@
As explained in Angius:2023uqk , the intersection of two EotW branes can be understood as providing a boundary to the theory living on each of the (possibly different) branes, or, more in line with the interpretation we will give now, as domain walls between the different boundaries of the theory.
@@
In the presence of $n$ compact dimensions, we can understand the intersection of an arbitrary number $k<d$ of EotW branes from the perspective of $k$-categories Atiyah:1989vu ; lurie2009 . From a categorical point of view, the bordisms on $n$-manifolds define a category $\mathbf{Cob}(n+1)$ in the following way:
@@
  • Composition of morphisms is given by gluing morphisms together by their intermediate boundary.
@@
  • The elements of $\textbf{Cob}_{2}(n+2)$ are closed oriented manifolds of dimensions $n$.
@@
    In other words, we consider bordisms between bordisms. It can be shown (see lurie2009 for more details and some nuances in the definitions) that the above structures are well-defined, and as such ${\rm Hom}_{\mathscr{C}}(\mathcal{B},\mathcal{B}^{\prime})\neq\emptyset$, so that two bordisms of the same manifolds are always cobordant.
@@
For our purposes, we will restrict $\textbf{Cob}_{2}(n+2)$ to consist in $\{\mathcal{C}_{n},\emptyset\}$, with the empty set seen as a $n$-dimensional manifold, such that $\mathscr{C}={\rm Map}_{\textbf{Cob}_{2}(n+2)}(\mathcal{C}_{n},\emptyset)$ are the possible bordisms from our compact manifold to nothing. Each of this will correspond to a different EotW brane/BoN construction. Finally, the elements in ${\rm Hom}_{\mathscr{C}}(\mathcal{B}_{n+1},\mathcal{B}^{\prime}_{n+1})$ will be the bordisms between this, interpolating between intersecting EotW branes. An “angular” coordinate $\theta$ parallel to the two branes and perpendicular to the radial one $\rho$ can act as a new Morse-Bott function over $\mathcal{X}_{n+2}\in{\rm Hom}_{\mathscr{C}}(\mathcal{B}_{n+1},\mathcal{B}^{\prime}_{n+1})$. Following an analogous procedure to that of section 3, one could bound the topology of $\mathcal{X}_{n+2}$, and from that, the topology changes in going from $\mathcal{B}_{n+1}$ to $\mathcal{B}^{\prime}_{n+1}$, which can be then understood as the intersecting EotW branes. As the necessary defects to kill topological/cobordism charges should be the same, on principle no additional defects would be needed to be located at these intersections/topology changes (a feature also observed in the universal scalings studied in Angius:2023uqk ), with $\mathcal{B}_{n+1}$ and $\mathcal{B}^{\prime}_{n+1}$ differing only on their arrangement. See Figure 11 for a simple example in a toroidal compactification. This analysis could on principle allow us to obtain the qualitative properties and topological structure of intersecting EotW branes, which so far have not been thoroughly explored, though explicit examples appear in Angius:2023uqk .
@@
To illustrate the above reasoning, we consider a network of EotW branes in the type IIB on $\Sigma_{g}$ from section 4.2. For simplicity, we take $\Sigma_{1}\simeq\mathbb{T}^{2}$. There are two possible bordisms that do not imply the creation of new cycles, namely (A) the shrinking of one cycle over its dual such that $\mathcal{V}\to 0$ and the torus goes to nothing in a single step, and (B) the pinching of a cycle, such that there is a topology change to $\mathbb{S}^{2}$, which later shrinks to a point. As discussed below (4.24), these two bordisms are associated to distinct critical exponents, $\delta_{\textbf{A}}=\sqrt{\frac{8}{11}}$ and $\delta_{\textbf{B}}=1$. If the original $\mathbb{T}^{2}$ is threaded by $F_{1}$ flux, then this domain wall (which in (A) ends spacetime) corresponds with a D7-brane wrapped along the dual 1-cycle. This is illustrated in Figure 11. The angle $\alpha>0$ between the different EotW branes should be set by the tension of the unknown object/defect that causes the Morse-Bott function to have different critical points as we move in the parallel direction $\theta$. Understanding its nature would allow us to compute the value of said angle.
@@
Seeing $\{{\rm Crit}(\rho)|_{\theta}\}_{\theta}$ as a set parameterized by $\theta$ (for the case shown in Figure 11 this is not even a manifold), the point $\theta_{\star}$ where the fiber changes topology correspond with the location of the EotW branes intersection.
@@
as we move in the $\theta$ direction. A single type of defect would be enough to explain the same angle $\alpha$ between the branes at each intersection. Identifying its nature and the value of $\alpha$ could allow us to put an upper bound on the number of possible EotW branes, as these could not span more than the celestial sphere. See Angius:2023uqk for other explicit computations of the critical exponents in EotW brane intersection networks for compactifications on Calabi-Yau threefolds.
The above reasoning can be extended to intersections of more than two EotW branes. For this we recursively define $\textbf{Cob}_{k}(n+k)$ (restricted to $\{\mathcal{C}_{n},\emptyset\}$), progressively trading spatial directions with coordinates (acting as Morse-Bott functions) in the higher-dimensional bordism. The end point of these sequence of $k$-categories is given by $\textbf{Cob}_{d-1}(d+n-1)$, where every spatial macroscopic direction is capped by an EotW brane, resulting in a (spatially) finite-volume universe. This might be of interest in the context of “Bubbles of something” and the “boundary proposal” Friedrich:2023tid ; Friedrich:2024aad , see also Hawking:1998bn ; Turok:1998he ; Garriga:1998ri ; Bousso:1998pk , where the universe is nucleated by the reverse process of a bubble of nothing, in such a way that the boundary of these gravitational instantons could be described by local patches of different topology. In any case, we will not comment more on this and leave it for a future work.
@@ 6 Conclusions and Outlook
[…]
@@
Furthermore, we are able to extend our results and interpretation to more involved settings such as collision of Bubbles of Nothing, where we identify the possible existence (or lack) of topological obstructions resulting in curvature singularities before the tip of the BoN come into contact. We also provide an interpretation of the intersection of EotW branes through higher bordisms, which might help in the understanding of the junction structure.
@@
Our results are only a first step, however. The new findings are mostly in the topological, more qualitative, side of the bordisms to nothing. While in section 4.2 we commented on how, for type IIB compactifications on $\Sigma_{g}$ each topology change lowered the value of the vacuum energy, in general one would like to also obtain the decay rates associated to them, in line with the examples in GarciaEtxebarria:2020xsr and the general considerations on decay rates/nucleation probabilities on bubbles of nothing/something from Friedrich:2023tid ; Friedrich:2024aad . Smooth bordisms are interesting precisely for this purpose, as their action can be computed directly with EFT ingredients, without the need to include UV contributions whose behavior might not be well understood, and also diverge from the EFT perspective. One could expect that the different types of topology changes could be studied separately, in such a way that by considering the different chains (3.1) $\mathcal{C}_{n}=\mathcal{C}_{n}^{(0)}\to\mathcal{C}_{n}^{(1)}\to\dots\to\mathcal{C}_{n}^{(m)}=\emptyset$, with the total decay rate given as a sum of blocks associated to each of the changes, in such a way that different bordisms could be sorted by their decay rate. This would be interesting for obvious phenomenological reasons, but goes well over the scope of the paper, and it is left for future work.
[…]
@@
Apart from topological constrains, we have not considered the possible dynamical restrictions to vacuum decay. It is possible that, in the process of breaking some of them, specific decay channels or bordism topologies are singled-out or excluded. We leave for a future project the better understanding of the relation between these dynamical constrains and processes through which the topology/geometry of $\mathcal{B}_{n+1}$ changes. On the other hand, as explained in Appendix B and Saltman:2004jh ; Adams:2005rb , topology changes consisting in the decay of a 1-cycle can be explained through tachyon condensation. One expects that tachyon condensation is also involved in decay of higher dimensional cycles, but the explicit mechanism does not directly translate from the 1-cycle case. The worldsheet interpretation of $n$-cycle closing is an interesting problem we would like to return to in the future.
@@
Other type of restrictions would be those in which a fibration structure is required along the complete bordism direction. For example, F-theory compactifications need to keep the elliptic fiber for all values of $\rho$, and thus $\mathcal{B}_{n+1}$, with topology changes in which said fiber disappears not being considered. A possible solution would be trying to extend the $\mathbb{T}^{2}\hookrightarrow\mathcal{C}_{n}\to\mathcal{D}_{n-2}$ fibration to $\mathbb{T}^{2}\hookrightarrow\mathcal{B}_{n+1}\to\hat{\mathcal{D}}_{n-1}$, in such a way that $\partial\hat{\mathcal{D}}_{n-1}=\mathcal{D}_{n+2}$ and the bordism fibration restricting to that of $\mathcal{C}_{n}$ along the boundary. How to properly perform this is beyond the scope of this paper.
@@
It has been conjectured Strominger:1996it ; chan2014strominger that Calabi-Yau threefolds always admit a $\mathbb{T}^{3}$ fibration, what motivated the authors in GarciaEtxebarria:2020xsr to propose that the BoN solutions constructed for spin bordisms on this manifold (see also Acharya:2019mcu for the additional bordisms of $\mathbb{T}^{3}$ to nothing) could be used as decay channels for realistic CY3 compactifications. However, as we have shown with examples such as precisely $\mathbb{T}^{2}\hookrightarrow\mathcal{B}_{4}\to\mathbb{D}^{2}$ bordism for $\mathbb{T}^{3}$ in section 4.3, bordisms given by non trivial fibrations have different homology groups than their trivial counterparts, resulting in a different topology change sequence. While the claim of GarciaEtxebarria:2020xsr probably holds in the sense that an analogous decay channel exists, it is not immediate that the constructions known in the literature can be straightforwardly used, as the $\mathbb{T}^{3}$ fibration might not be trivial.
@@
Although we have simply considered the $\rho$ component of the bordism metric as our Morse function, it is easy to see that any $p$-form field $\omega$ defined on $\mathcal{C}_{n}$ can be extended into the whole bordism $\mathcal{B}_{n+1}$ as a $(p+1)$-form by taking $\omega^{\prime}=\omega\wedge\mathrm{d}\rho$, being zero at critical points of $\rho$. It might be interesting to relate general properties of these fields with those of the bordism. Furthermore, we would like to point out that only (co)homology coefficients such as Betti numbers and torsion ranks have been considered. We leave for future work the possibility of recovering other topological invariants of the bordism from those of $\mathcal{C}_{n}$, as well as understanding to which dynamical properties of the bordism to nothing they are related to.
@@
There is still much to be understood and computed about the possible channels through which our universe can decay to nothing. Computations are usually quite involved, so that apart from specific cases GarciaEtxebarria:2020xsr , most constructions in the literature feature relatively simple compactifications. We believe that studying more complicated cases can lead to a better understanding of how our universe can be created and disappear into nothing is within reach, and we expect this paper serves as a small step in this direction.

### v2

@@ 5 Stories with more than one ending
In the above sections, only a single BoN/EotW brane was considered for each case. However, instances of multi-centered BoN solutions eventually colliding or networks of EotW branes intersecting are the natural generalization. These cases have been previously studied in the literature (see Horowitz and Maeda (2002); Freivogel et al. (2007); Kleban (2011); Moritz and Valenzuela (2024) and Angius et al. (2024a); Angius (2024); García Etxebarria et al. (2024), respectively), and as we will see, we can learn important qualitative properties through our approach.
@@ 5.1 Seeing double: Collisions of BoN’s
As explained in section 2.1, bubbles of nothing have some non-zero probability of nucleating per unit volume. It is then expected, that, in a large macroscopic region and for a long enough period of time, several BoN nucleate and start growing, eventually approaching each other’s surfaces. Considering our lower-dimensional theory to result from compactification on a compact surfaces $\mathcal{C}_{n}$, a bubble solution will be associated to a bordism and Morse function choice, $(\mathcal{B}_{n+1},\rho)$. Two different BoN can in principle be associated to $\mathcal{B}_{n+1}$ and $\mathcal{B}_{n+1}^{\prime}$, such that $\mathcal{C}_{n}=\partial\mathcal{B}_{n+1}=\partial\mathcal{B}_{n+1}^{\prime}$, without necessarily the two bordisms being homeomorphic. Moving then radially from the surface of one BoN to the other, this can be seen as traveling along some compact $(n+1)$-manifold $\mathcal{X}_{n+1}$ given by
@@
As the surfaces of the two bubbles approach, the volume of $\mathcal{X}_{n+1}$ shrinks. Whilst the eventual collision of the two tips can reach Planckian curvatures, as well as being affected by the backreaction of large matter densities being dragged by the bubble, one can also have these problems well before the BoN surface if $\mathcal{X}_{n}$ cannot be contracted to a point in a single topology change, with the curvatures of internal cycles of $\mathcal{X}_{n+1}$ blowing up before that of the whole manifold. Having a single topology change translates in the Morse function interpolating between the tips of the two bubbles having only critical two points, respectively with index 0 and $n+1$, associated with the tip of each BoN. From (3.14), this translates in
@@
which implies (splitting between free and torsion parts, see footnote 8)
@@
In the absence of torsion (i.e., all $t_{k}(\mathcal{M})\equiv 0$), a homology sphere solution $\vec{b}(\mathcal{X}_{n+1})=\vec{b}(\mathbb{S}^{n+1})$ is compatible [^24] Note that this does not imply such solution exists, but rather that there is no immediate obstruction to its existence. with the above (5.2) and (5.4) equations, by choosing [^25] For this choice, the resulting ${\mathrm{D}}\mathcal{B}_{n+1}\simeq\left(\mathcal{B}_{n+1}\cup\overline{\mathcal{B}_{n+1}}\right)/_{\sim_{\partial\mathcal{B}_{n+1}}}$ is called double of $\mathcal{B}_{n+1}$. $\mathcal{B}_{n+1}^{\prime}\simeq\mathcal{B}_{n+1}$ and saturating (3.8). This is illustrated for $\mathcal{C}_{2}\simeq\Sigma_{2}$ in Figure 10. This means that in principle there should not be a topological obstruction to the collision of two BoN with the compact manifold $\mathcal{X}_{n+1}$ between them shrinking to a point.
@@
As a final comment, we note that, unlike in the case of a single BoN, which required the presence of D(8p)(8-p)-branes as cobordisms defects when considering $p$-form fluxes, here their presence is much more mild than the intermediate topology changes commented above. D(8p)(8-p)-branes, located along along $\mathcal{B}_{n+1}$ and $\mathcal{B}^{\prime}_{n+1}$, needed to kill the $p$-form charges, have opposite orientations, and thus will annihilate each other when they collide as $\mathcal{X}_{n+1}$ closes. Another way to see this is by considering that the (co)homology groups of $\mathcal{X}_{n+1}$ and $\mathcal{C}_{n}$ are different, and as commented above, $H_{k}(\mathcal{X}_{n+1},\mathbb{Z})\simeq H^{k}(\mathcal{X}_{n+1},\mathbb{Z})$ can be trivial for $k=1,\,\dots,\,n$ in the absence of torsion, meaning that the number of $p$-form fields/charges is no longer conserved on the compact manifold connecting the two BoN.
Going back to the non-minimal $\mathbb{T}^{3}$ bordism in section 4.3, as explained in García Etxebarria et al. (2020), $\partial\mathcal{B}_{4}$ can be thought as “half a K3” manifold, understood as an elliptical fibration over a $\mathbb{P}^{1}$ base, with 24 degenerations. This way, the double of our bordism is indeed ${\rm D}\mathcal{B}_{4}={\rm K3}$, which is not a homology 4-sphere, and as such, cannot be shrunk to a point without intermediate topology changes. The collision of two BoN such as those described in García Etxebarria et al. (2020) would require Planckian curvatures and stringy effects, even if these were not needed for the single BoN construction.
@@ 5.2 Going a bordism further: Intersection of EotW branes
We finally consider the cases with more than one bordism to nothing, in the language of EotW branes. Intersection of these was studied in Angius et al. (2024a); Angius (2024); García Etxebarria et al. (2024), and in this section we will study them from the point of view of topology changes between “cobordant” bordisms to nothing.
From the EFT point of view, the single EotW brane from section 2.2 was generalized to $k$ intersecting EotW branes in Angius et al. (2024a). For this the following lower-$d$ EFT action is considered,
@@
where the moduli space metric has unit diagonal, $\mathsf{G}_{ii}=1$, as well as the following ansatz for a conformally flat metric (see Angius (2024) for EotW solutions with non-conformally flat backgrounds):
@@
As explained in Angius et al. (2024a), the intersection of two EotW branes can be understood as providing a boundary to the theory living on each of the (possibly different) branes, or, more in line with the interpretation we will give now, as domain walls between the different boundaries of the theory.
@@
In the presence of $n$ compact dimensions, we can understand the intersection of an arbitrary number $k<d$ of EotW branes from the perspective of $k$-categories Atiyah (1989); Lurie (2009). From a categorical point of view, the bordisms on $n$-manifolds define a category $\mathbf{Cob}(n+1)$ in the following way:
@@
  • Composition of morphisms is given by gluing bordisms together by their intermediate boundary.
@@
  • The elements of $\textbf{Cob}_{2}(n+2)$ are closed oriented manifolds of dimension $n$.
@@
    In other words, we consider bordisms between bordisms. It can be shown (see Lurie (2009) for more details and some nuances in the definitions) that the above structures are well-defined, and as such ${\rm Hom}_{\mathscr{C}}(\mathcal{B},\mathcal{B}^{\prime})\neq\emptyset$, so that two bordisms of the same manifolds are always cobordant.
@@
For our purposes, we will restrict $\textbf{Cob}_{2}(n+2)$ to consist in $\{\mathcal{C}_{n},\emptyset\}$, with the empty set seen as a $n$-dimensional manifold, such that $\mathscr{C}={\rm Map}_{\textbf{Cob}_{2}(n+2)}(\mathcal{C}_{n},\emptyset)$ are the possible bordisms from our compact manifold to nothing. Each of this will correspond to a different EotW brane/BoN construction. Finally, the elements in ${\rm Hom}_{\mathscr{C}}(\mathcal{B}_{n+1},\mathcal{B}^{\prime}_{n+1})$ will be the bordisms between this, interpolating between intersecting EotW branes. An “angular” coordinate $\theta$ parallel to the two branes and perpendicular to the radial one $\rho$ can act as a new Morse-Bott function over $\mathcal{X}_{n+2}\in{\rm Hom}_{\mathscr{C}}(\mathcal{B}_{n+1},\mathcal{B}^{\prime}_{n+1})$. Following an analogous procedure to that of section 3, one could bound the topology of $\mathcal{X}_{n+2}$, and from that, the topology changes in going from $\mathcal{B}_{n+1}$ to $\mathcal{B}^{\prime}_{n+1}$, which can be then understood as the intersecting EotW branes. As the necessary defects to kill topological/cobordism charges should be the same, in principle no additional defects would be needed to be located at these intersections/topology changes (a feature also observed in the universal scalings studied in Angius et al. (2024a)), with $\mathcal{B}_{n+1}$ and $\mathcal{B}^{\prime}_{n+1}$ differing only on their arrangement. See Figure 11 for a simple example in a toroidal compactification. This analysis could in principle allow us to obtain the qualitative properties and topological structure of intersecting EotW branes, which so far have not been thoroughly explored, though explicit examples appear in Angius et al. (2024a).
@@
To illustrate the above reasoning, we consider a network of EotW branes in the type IIB on $\Sigma_{g}$ from section 4.2. For simplicity, we take $\Sigma_{1}\simeq\mathbb{T}^{2}$. There are two possible bordisms that do not imply the creation of new cycles, namely (A) the shrinking of one cycle over its dual such that $\mathcal{V}\to 0$ and the torus goes to nothing in a single step, and (B) the pinching of a cycle, such that there is a topology change to $\mathbb{S}^{2}$, which later shrinks to a point. As discussed below (4.24), these two bordisms are associated to distinct critical exponents, $\delta_{\textbf{A}}=\sqrt{\frac{8}{11}}$ and $\delta_{\textbf{B}}=1$. If the original $\mathbb{T}^{2}$ is threaded by $F_{1}$ flux, then this domain wall (which in (A) ends spacetime) corresponds to a D7-brane wrapped along the dual 1-cycle. This is illustrated in Figure 11. The angle $\alpha>0$ between the different EotW branes should be set by the tension of the unknown object/defect that causes the Morse-Bott function to have different critical points as we move in the parallel direction $\theta$. Understanding its nature would allow us to compute the value of said angle.
@@
Seeing $\{{\rm Crit}(\rho)|_{\theta}\}_{\theta}$ as a set parameterized by $\theta$ (for the case shown in Figure 11 this is not even a manifold), the point $\theta_{\star}$ where the fiber changes topology correspond to the location of the EotW branes intersection.
@@
as we move in the $\theta$ direction. A single type of defect would be enough to explain the same angle $\alpha$ between the branes at each intersection. Identifying its nature and the value of $\alpha$ could allow us to put an upper bound on the number of possible EotW branes, as these could not span more than the celestial sphere. See Angius et al. (2024a) for other explicit computations of the critical exponents in EotW brane intersection networks for compactifications on Calabi-Yau threefolds.
The above reasoning can be extended to intersections of more than two EotW branes. For this we recursively define $\textbf{Cob}_{k}(n+k)$ (restricted to $\{\mathcal{C}_{n},\emptyset\}$), progressively trading spatial directions with coordinates (acting as Morse-Bott functions) in the higher-dimensional bordism. The end point of these sequence of $k$-categories is given by $\textbf{Cob}_{d-1}(d+n-1)$, where every spatial macroscopic direction is capped by an EotW brane, resulting in a (spatially) finite-volume universe. This might be of interest in the context of “Bubbles of something” and the “boundary proposal” Friedrich et al. (2024); Friedrich and Hebecker (2024), see also Hawking and Turok (1998); Turok and Hawking (1998); Garriga (1998); Bousso and Chamblin (1999), where the universe is nucleated by the reverse process of a bubble of nothing, in such a way that the boundary of these gravitational instantons could be described by local patches of different topology. In any case, we will not comment more on this and leave it for a future work.
@@ 6 Conclusions and Outlook
In this paper, we have given a new interpretation of spacetime bordisms to nothing through Morse-Bott theory. The radial distance to the end of spacetime is interpreted as a Morse function over the bordism, and its fibers or level curves correspond to the internal manifold at each spacetime point. We are thus able to obtain properties of the internal topology changes for bordisms to nothing admitting a topological manifold description, learning about which of these changes must always occur, and where defects, such as Dpp-branes, must be located. Though qualitative in nature, these results serve as a first step towards a piece-wise construction of spacetime ending configurations, such as BoN solutions, for more involved internal geometries. Knowing their topology, it might be possible to proceed analogously to García Etxebarria et al. (2020), dividing $\mathcal{B}_{n+1}$ into different regions whose metric can be given an analytic approximate expression, and later gluing them together. As happens with the analysis in section 4.2, for large volume regimes in $\mathcal{C}_{n}$ it is consistent to expect that topology changes on the compact manifold, consisting in the shrinking of some internal cycles, do not immediately trigger topology changes for the rest of cycles. This way it is possible to have an EFT description between the different domain walls.
@@
Furthermore, we are able to extend our results and interpretation to more involved settings such as collision of Bubbles of Nothing, where we identify the possible existence (or lack) of topological obstructions resulting in curvature singularities before the tips of the BoNs come into contact. We also provide an interpretation of the intersection of EotW branes through higher bordisms, which might help in the understanding of the junction structure.
@@
Our results are only a first step, however. The new findings are mostly in the topological, more qualitative, side of the bordisms to nothing. While in section 4.2 we commented on how, for type IIB compactifications on $\Sigma_{g}$ each topology change lowered the value of the vacuum energy, in general one would like to also obtain the decay rates associated to them, in line with the examples in García Etxebarria et al. (2020), and the general considerations on decay rates/nucleation probabilities on bubbles of nothing/something from Friedrich et al. (2024); Friedrich and Hebecker (2024). Smooth bordisms are interesting precisely for this purpose, as their action can be computed directly with EFT ingredients, without the need to include UV contributions whose behavior might not be well understood, and also can diverge from the EFT perspective. One could expect that the different types of topology changes could be studied separately, in such a way that by considering the different chains (3.1) $\mathcal{C}_{n}=\mathcal{C}_{n}^{(0)}\to\mathcal{C}_{n}^{(1)}\to\dots\to\mathcal{C}_{n}^{(m)}=\emptyset$, with the total decay rate given as a sum of blocks associated to each of the changes, in such a way that different bordisms could be sorted by their decay rate. This would be interesting for obvious phenomenological reasons, but goes well over the scope of the paper, and it is left for future work.
Another matter is that of the defects considered. As commented above, for our approach to apply, we need to have smooth bordisms, or at most mild defects such as conical singularities which can be “smoothed out”. While cobordism defects such as Dpp-branes fulfill this, and indeed we have been able to infer their arrangement along the bordism, other more singular defects, such as O-planes, or more exotic branes McNamara and Vafa (2019); Dierigl et al. (2024), might not guarantee that. It would be nice to understand which types of defects allow for some smooth(able) manifold description and which do not. As evidenced by the example from Dierigl et al. (2024) shown in Figure 5(b), some of these defects or configurations require for non-trivial cycles along $\mathcal{B}_{n+1}$ (resulting in thick domain walls), and thus cannot be located along minimal bordisms. A general understanding of when more involved bordisms are required would also have important implications for the construction of realistic BoN.
@@
Apart from topological constrains, we have not considered the possible dynamical restrictions to vacuum decay. It is possible that, in the process of breaking some of them, specific decay channels or bordism topologies are singled-out or excluded. We leave for a future project the better understanding of the relation between these dynamical constrains and processes through which the topology/geometry of $\mathcal{B}_{n+1}$ changes. On the other hand, as explained in Appendix B and Saltman and Silverstein (2006); Adams et al. (2005), topology changes consisting in the decay of a 1-cycle can be explained through tachyon condensation. One expects that tachyon condensation is also involved in decay of higher dimensional cycles, but the explicit mechanism does not directly translate from the 1-cycle case. The worldsheet interpretation of $n$-cycle closing is an interesting problem we would like to return to in the future.
@@
Other type of restrictions would be those in which a fibration structure is required along the complete bordism direction. For example, F-theory compactifications need to keep the elliptic fiber for all values of $\rho$, and thus $\mathcal{B}_{n+1}$, with topology changes in which said fiber disappears not being considered. A possible solution would be trying to extend the $\mathbb{T}^{2}\hookrightarrow\mathcal{C}_{n}\to\mathcal{D}_{n-2}$ fibration to $\mathbb{T}^{2}\hookrightarrow\mathcal{B}_{n+1}\to\hat{\mathcal{D}}_{n-1}$, in such a way that $\partial\hat{\mathcal{D}}_{n-1}=\mathcal{D}_{n-2}$ and the bordism fibration restricting to that of $\mathcal{C}_{n}$ along the boundary. How to properly perform this is beyond the scope of this paper.
@@
It has been conjectured Strominger et al. (1996); Chan (2014) that Calabi-Yau threefolds always admit a $\mathbb{T}^{3}$ fibration, what motivated the authors in García Etxebarria et al. (2020) to propose that the BoN solutions constructed for spin bordisms on this manifold (see also Acharya (2020) for the additional bordisms of $\mathbb{T}^{3}$ to nothing) could be used as decay channels for realistic CY3 compactifications. However, as we have shown with examples such as precisely $\mathbb{T}^{2}\hookrightarrow\mathcal{B}_{4}\to\mathbb{D}^{2}$ bordism for $\mathbb{T}^{3}$ in section 4.3, bordisms given by non trivial fibrations have different homology groups than their trivial counterparts, resulting in a different topology change sequence. While the claim of García Etxebarria et al. (2020) probably holds in the sense that an analogous decay channel exists, it is not immediate that the constructions known in the literature can be straightforwardly used, as the $\mathbb{T}^{3}$ fibration might not be trivial.
@@
Although we have simply considered the $\rho$ component of the bordism metric as our Morse function, it is easy to see that any $p$-form field $\omega$ defined on $\mathcal{C}_{n}$ can be extended into the whole bordism $\mathcal{B}_{n+1}$ as a $(p+1)$-form by taking $\omega^{\prime}=\omega\wedge\mathrm{d}\rho$, being zero at critical points of $\rho$. It might be interesting to relate general properties of these fields with those of the bordism. Furthermore, we would like to point out that only (co)homology coefficients such as Betti numbers and torsion ranks have been considered. We leave for future work the possibility of recovering other topological invariants of the bordism from those of $\mathcal{C}_{n}$, as well as understanding to which dynamical properties of the bordism to nothing they are related to. A first step for this would be the derivation of generalizations to the Morse-Bott inequalities (3.12) which also take into account, e.g., the spin structure along the boundary, not only its homology. This could result in higher lower bounds on the number of topology changes, maybe finding that the 13 topology changes for the $\mathbb{T}^{3}$ with periodic boundary conditions from section 4.3 saturate said new bounds.
@@
There is still much to be understood and computed about the possible channels through which our universe can decay to nothing. Computations are usually quite involved, so that apart from specific cases García Etxebarria et al. (2020), most constructions in the literature feature relatively simple compactifications. We believe that studying more complicated cases, leading to a better understanding of how our universe can be created and disappear into nothing, is within reach, and we expect that this paper serves as a small step in this direction.